OpenStudy (anonymous):

Grace starts to save at age 20 for an extended vacation around the world that she will take on her 45th birthday. She will contribute $250 four times each year to the account, which earns 1.85% annual interest, compounded annually. What is the future value of this investment when she takes her trip?

OpenStudy (anonymous):

hi if the account has daily return calculation : fv = 250 * (((1+1.85 % / 4)^ 25*4) - 1 ) / (1.85% /4) fv = 250 * 126.778 = 31,694 $ but if account compounds at the end of year : fv = 1,000 * (((1+ 1.85%)^25) - 1) / 1.85% fv= 1,000 * 31.423 = 31,423 $ hope it helps

OpenStudy (anonymous):

thanks!